OpenStudy (anonymous):

If cosx=2/3 and x is in quadrant 4: Find sinx,tanx, and sec x

OpenStudy (anonymous):

cos x=base/hypotenuse compare with given info.. base=2 hypotenuse=3 also cosx=+ve in 4th quadrant now solve for sinx tanx and sec x.. find perependicular using pythogoras then sinx=perp/hyp tanx=p/b sec x=1/cosx

OpenStudy (anonymous):

Also ,sin x , tan x are -ve in 4th quadrant